OpenStudy (anonymous):

An Indian tepee is a conical tent. Find the number of cubic feet of air in a teepee 10 ft. in diameter and 12 ft. high. the answer has to be like this ____pii cu.ft

OpenStudy (shubhamsrg):

use this -> volume of cone = 1/3 pi (radius^2)(height) just insert values..radius=diameter/2

OpenStudy (anonymous):

i tried that.. i got 13 :/

OpenStudy (anonymous):

is it 25/3?

OpenStudy (anonymous):

answerr me :(

OpenStudy (shubhamsrg):

sorry for late reply : its 1/3(5^2)(12)(pi) = 100pi

OpenStudy (anonymous):

really?

OpenStudy (anonymous):

howwwwwwwwwwww?

OpenStudy (shubhamsrg):

lol..yes..thats the bitter truth :P

OpenStudy (anonymous):

wait i did that!

OpenStudy (shubhamsrg):

well radius of cone = diameter/2 = 10/2 = 5 and volume = 1/3(pi)(r^2)(h)

OpenStudy (anonymous):

then....

OpenStudy (shubhamsrg):

just substitute values mam,,r=5,,h=12 ..and you get 100pi

OpenStudy (anonymous):

omg i was like multiplyin 25 and 12 thenn dividing it by 3..

OpenStudy (shubhamsrg):

yes,,thats what has to be done mam,,you'll get 100 only like this..

OpenStudy (anonymous):

thank you very much!

OpenStudy (shubhamsrg):

never mind ..glad to help ^^

OpenStudy (anonymous):

could u help me wih this one... please An Indian tepee is a conical tent. Find the number of skins needed to cover a teepee 10 ft. in diameter and 12 ft. high. Each skin covers 15 sq. ft. (use = 3.14)

OpenStudy (shubhamsrg):

well this time it asks you for the curved surface area of the cone..calculate it and divide by 15..you'll get your answer..

OpenStudy (anonymous):

so i just have to divide 100 by 15 right??

OpenStudy (shubhamsrg):

curved surface area = 2(pi)(r)(l) where l = (h^2 + r^2)^(1/2)

OpenStudy (shubhamsrg):

100 was the volume,,this is a different part,,you need to calculate Curved surface area separately..

OpenStudy (anonymous):

ohh ok

OpenStudy (anonymous):

is it 408.2

OpenStudy (anonymous):

i used surface area = 2(pi)(r)(l)

OpenStudy (shubhamsrg):

surface area i guess you have calculated correctly.. was your l= 13 ?

OpenStudy (anonymous):

yup:)

OpenStudy (anonymous):

you rock! thanks:)

OpenStudy (shubhamsrg):

then i guess its correct..just divide by 15 and you get your answer..

OpenStudy (shubhamsrg):

hehe..nevermind,, ^_^

OpenStudy (anonymous):

is it 27?

OpenStudy (shubhamsrg):

approximately yes,,otherwise there is a fraction here and there..

OpenStudy (anonymous):

lol okay c:

OpenStudy (shubhamsrg):

to be more precise,,its around 27.213...

OpenStudy (anonymous):

oh okay:)

OpenStudy (shubhamsrg):

mam am so very sorry :( CSA = pi(r)(l) and not 2pi(r)(l)

OpenStudy (shubhamsrg):

but you dont need to calculate from the starting,,your final answer is juts half of what you got previously..

OpenStudy (shubhamsrg):

am really sorry :/